40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 4 10:51:10 2012, Steve, Update, SUS, PRM damping restored 
    Reply  Mon Dec 17 14:13:55 2012, Jenne, Update, SUS, PRM oplev gains restored 
Message ID: 7781     Entry time: Tue Dec 4 10:51:10 2012     Reply to this: 7838
Author: Steve 
Type: Update 
Category: SUS 
Subject: PRM damping restored 

PRM oplev gains set to zero from PIT 0.15 and YAW -0.3 and damping restored

ELOG V3.1.3-